# Video-3D LLM: Learning Position-Aware Video Representation for 3D Scene Understanding

### v1

@@ 4.2.1 Comparison Baselines
2D LLMs. Oryx [35] has included the ScanQA dataset in its training stage. In addition, we test the zero-shot performance of LLaVA-Video [53].
@@ 4.2.2 Results
3D Visual Grounding. For 3D visual grounding, we follow previous work [24, 58, 22] to detect all objects using Mask3D [39] and then make predictions over the object proposals. Our model achieves the highest accuracy, with Acc@0.25 at 58.1% and Acc@0.5 at 51.7% on ScanRefer, and F1@0.25 at 58.0% and F1@0.5 at 52.7% on Multi3DRefer. As previous 3D LLMs either use detected object proposals (e.g., ChatScene, Chat3D) or train an additional grounding module based on a 3D detector (e.g., Grounded 3d-LLM, PQ3D, LLaVA-3D), we can compare with these 3D LLMs fairly. Specifically, our model improves Acc@0.25 by 2.6% on ScanRefer and F1@0.25 by 0.9% on Multi3DRefer compared to ChatScene, which uses the same object proposal as ours.
3D Dense Captioning. Following the previous setting [23, 22], we generate captions for each detected object proposal. Our method demonstrates superior performance in Scan2Cap, achieving 41.3 at B-4@0.5 and 83.8 at C@0.5. The results reveal that our method connects video content with its position information by injecting 3D-PE into video representations.
3D Question Answering. Our model outperforms best competitors with 30.1% EM on ScanQA and 58.6% EM on SQA3D, which could be attributed to the strong representations inherited from Video LLMs. It is worth noting that existing Video LLMs achieve competent results in a zero-shot manner, suggesting that current 3D question answering tasks may not sufficiently address the challenges of spatial understanding in 3D scenes.
3D-PE Coord. ScanRefer Scan2Cap ScanQA
Acc@0.25 Acc@0.5 C@0.5 EM
None Avg 57.50 50.84 31.03 30.03
MLP 59.63 52.98 76.23 29.62
Sin 58.11 51.72 83.76 30.09
Sin Center 57.53 51.06 80.88 29.39
Min-Max 58.05 51.77 82.75 30.18
Avg 58.11 51.72 83.76 30.09
@@
4.3 Ablation Study
Effectiveness of Frame Sampling. In Table 2, we assess the effectiveness of the frame sampling strategy across varying numbers of frames. To evaluate the model’s efficiency, we calculated the average inference speed on ScanQA. In the fixed frame number setting, we sample frames until reaching the desired number. As shown in the table, the model’s performance improves with an increasing number of frames, though inference time also rises. Video-3D LLM surpasses many previous methods with only 8 frames, demonstrating that our position-aware video paradigm can effectively model 3D scene understanding tasks. With 8 and 16 frames, maximum coverage sampling (“MC”) significantly enhances performance for all tasks by capturing a more complete 3D scene. Notably, with 8 frames, it improves Acc@0.25 by $4.54$ on ScanRefer and C@0.5 by $4.26$ on Scan2Cap compared to the uniform sampling strategy (“Uniform”). With 32 frames, the “Uniform” strategy shows comparable results with “MC”, as this number of frames is more than enough to cover the 3D scene. In the adaptive frame number setting, frames are sampled until over 95% of the scene’s voxels are covered or a maximum of 32 frames is reached. This allows for flexible adjustment based on scene size. As shown in the results, “MC” uses an average of $18$ frames across all scenes, achieving similar performance to the 32-frame uniform strategy while offering better inference speed. Meanwhile, compared to LLaVA-3D [57], “MC” shows superior performance while achieving similar inference speed.
@@ 4.3 Ablation Study
Effectiveness of Position-Aware Video Representation. Table 3 presents the performance using different 3D position encoding (3D-PE) and coordinate aggregation strategies. The model was trained on the entire multi-task dataset, with 32 frames under the uniform frame sampling. We first assess the impact of different 3D-PE using average coordinate aggregation. The results indicate that the model can perform well without 3D-PE on 3D grounding and 3D QA tasks. This suggests that the video LLM paradigm can effectively understand spatial relationships across multiple frames. However, for dense captioning tasks, which require locating objects within video sequences based on query bounding boxes, the absence of 3D-PE results in significantly degraded performance. Additionally, MLP position encoding proves more effective for grounding tasks, while Sin PE works better for others. We then evaluate different coordinate aggregation strategies using Sin position encoding. Since the 3D coordinates are more complex and variable than those in 2D images, simply using the coordinates at the center of a 2D patch doesn’t accurately reflect the spatial location. As shown in the result, using average 3D coordinates better represents the 3D location within a patch, while using the minimum and maximum 3D coordinates within the patch is also a viable alternative.
@@ 5 Conclusion
In this paper, we propose a generalist model for 3D scene understanding, namely Video-3D LLM. It is built upon a Video LLM and incorporates 3D coordinates into video representations. By doing this, it aligns video representations with 3D physical world, thereby enabling a suite of 3D scene understanding tasks. In addition, we introduce a maximum coverage sampling strategy to balance computational costs with performance efficiency. Our extensive experimental results demonstrate that our model achieves state-of-the-art performance on several benchmarks.
@@ Appendix A Implementation Details
Dataset Statistics. We present the detailed statistics for training and testing data in Table 5 and 6, respectively. Following previous work [23, 22], we adopt the validation set for ScanRefer [6], Multi3DRefer [52], Scan2Cap [7], ScanQA [7], and the test set for SQA3D [36]. All data have been converted to LLaVA [34] format, and we conduct statistics in this format.
@@ Appendix B Detailed Comparison
Scan2Cap. As shown in Table 8, we provide a detailed comparison on the validation set of Scan2Cap [7]. During inference, we utilize the object proposals from [23], which include 50 predicted objects extracted with Mask3D [39] for each scan. From the table, we can see our method achieves the state-of-the-art on CIDEr and BLEU-4 at 83.77 and 42.43, respectively.

### v2

@@ 4.2.1 Comparison Baselines
2D LLMs. Oryx [30] has included the ScanQA dataset in its training stage. We also test the zero-shot performance of LLaVA-Video [47].
@@ 4.2.2 Results
3D Visual Grounding. For 3D visual grounding, we follow previous work [19, 52, 17] to detect all objects and then make predictions over the object proposals [^2] For a fair comparison, we use the Mask3D-generated object proposals provided by LEO [18] for both 3D visual grounding and dense captioning.. Our model achieves the highest accuracy, with Acc@0.25 at 58.1% and Acc@0.5 at 51.7% on ScanRefer, and F1@0.25 at 58.0% and F1@0.5 at 52.7% on Multi3DRefer. As previous 3D LLMs either use detected object proposals (e.g., ChatScene, Chat3D) or train an additional grounding module based on a 3D detector (e.g., Grounded 3d-LLM, PQ3D, LLaVA-3D), we can compare with these 3D LLMs fairly. Specifically, our model improves Acc@0.25 by 2.6% on ScanRefer and F1@0.25 by 0.9% on Multi3DRefer compared to ChatScene, which uses the same object proposal as ours.
3D Dense Captioning. Following the previous setting [18, 17], we generate captions for each detected object proposal2. Our method demonstrates superior performance in Scan2Cap, achieving 41.3 at B-4@0.5 and 83.8 at C@0.5. The results reveal that our method connects video content with its position information by injecting 3D-PE into video representations.
3D Question Answering. Our model outperforms the best competitors with 30.1% EM on ScanQA and 58.6% EM on SQA3D, which could be attributed to the strong representations inherited from Video LLMs. Existing Video LLMs achieve competent results in a zero-shot manner, suggesting that current 3D QA tasks may not sufficiently address the challenges of spatial reasoning in 3D scenes.
4.3 Ablation Study
3D-PE Coord. Scan2Cap ScanRefer ScanQA
C@0.5 Acc@0.25 Acc@0.5 EM
None Avg 31.03 57.50 50.84 30.03
MLP 76.23 59.63 52.98 29.62
Sin 83.76 58.11 51.72 30.09
Sin Center 80.88 57.53 51.06 29.39
Min-Max 82.75 58.05 51.77 30.18
Avg 83.76 58.11 51.72 30.09
@@ 4.3 Ablation Study
Effectiveness of Frame Sampling. In Table 2, we assess the effectiveness of the frame sampling strategy across varying numbers of frames. To evaluate the model’s efficiency, we calculated the average inference speed on ScanQA. In the fixed frame number setting, we sample frames until reaching the desired number. As shown in the table, the model’s performance improves with an increasing number of frames, though inference time also rises. Video-3D LLM surpasses many previous methods with only 8 frames, demonstrating that our position-aware video paradigm can effectively model 3D scene understanding tasks. With 8 and 16 frames, maximum coverage sampling (“MC”) significantly enhances performance for all tasks by capturing a more complete 3D scene. Notably, with 8 frames, it improves Acc@0.25 by $4.54$ on ScanRefer and C@0.5 by $4.26$ on Scan2Cap compared to the uniform sampling strategy (“Uniform”). With 32 frames, the “Uniform” strategy shows comparable results with “MC”, as this number of frames is more than enough to cover the 3D scene. In the adaptive frame number setting, frames are sampled until over 95% of the scene’s voxels are covered or a maximum of 32 frames is reached. This allows for flexible adjustment based on scene size. As shown in the results, “MC” uses an average of $18$ frames across all scenes, achieving similar performance to the 32-frame uniform strategy while offering better inference speed. Meanwhile, compared to LLaVA-3D [51], “MC” shows superior performance while achieving similar inference speed. Maximum coverage sampling is performed once per scene, and the average time spent on ScanQA is only 17.8ms per question, which is negligible compared to the inference time.
@@
Scan2Cap ScanRefer ScanQA
C@0.5 B@0.5 Acc@0.25 Acc@0.5 EM
Voxel 54.9 34.3 56.1 49.8 28.1
Video 83.8 (+28.9) 41.3 (+7.0) 58.1 (+2.0) 51.7 (+1.9) 30.1 (+2.0)
Table 5: Ablation study for the 3D scene representation.
@@
Effectiveness of 3D-PE. Table 3 presents the performance using different 3D position encoding (3D-PE) and coordinate aggregation strategies. The model was trained on the entire multi-task dataset, with 32 frames under the uniform frame sampling. We first assess the impact of different 3D-PE using average coordinate aggregation. The introduction of 3D-PE leads to a consistent performance increase in 3D dense captioning and 3D visual grounding in all aggregation variants. Specifically, for dense captioning tasks, which require locating objects within video sequences based on query bounding boxes, the absence of 3D-PE results in significantly degraded performance. Additionally, MLP position encoding proves more effective for grounding tasks, while Sin PE works better for others. We then evaluate different coordinate aggregation strategies using Sin position encoding. Since the 3D coordinates are more complex and variable than those in 2D images, simply using the coordinates at the center of a 2D patch doesn’t accurately reflect the spatial location. As shown in the result, using average 3D coordinates better represents the 3D location within a patch, while using the minimum and maximum 3D coordinates within the patch is also a viable alternative.
@@
Effectiveness of Representing 3D Scenes as Videos. To eliminate the influence of the backbone, we conducted ablation experiments by comparing our 3D-as-video paradigm with voxel modeling based on the same backbone, LLaVA-Video. Following LLaVA-3D, we average the patch features corresponding to the same voxel to obtain the voxel features, and sample 3,096 of these features as input for the LLM. As shown in Table 5, there is a notable improvement of 28.9 C@0.5 on Scan2Cap, from 83.8 to 54.9. Note that our dataset is only 26% of LLaVA-3D, resulting the lower results compared to the LLaVA-3D. For ScanQA, which relies more heavily on visual perception, our approach still achieves a 2% gain in accuracy compared to voxel modeling. This indicates that our approach leverages spatiotemporal priors specifically, rather than merely benefiting from stronger visual features.
@@ 5 Conclusion
In this paper, we propose a novel paradigm to effectively exploit Video LLMs for 3D scene understanding. It incorporates 3D position encoding into video representations and employs a carefully-designed multi-task training recipe, thereby facilitating a suite of 3D scene understanding tasks. We further introduce a maximum coverage sampling strategy to optimize the trade-off between computational costs and model performance. Our extensive experimental results demonstrate the superiority of our method.
Acknowledgements
This work is supported by National Key R&D Program of China (Project No. 2022ZD0161200, 2022ZD0161201). This work is also supported by Hong Kong Research Grant Council - Early Career Scheme (Grant No. 24200223).
@@ Appendix A Implementation Details
Dataset Statistics. We present the detailed statistics for training and testing data in Table 6 and 7, respectively. Following previous work [18, 17], we adopt the validation set for ScanRefer [5], Multi3DRefer [46], Scan2Cap [6], ScanQA [6], and the test set for SQA3D [31]. All data have been converted to LLaVA [29] format, and we conduct statistics in this format.
@@ Appendix B Detailed Comparison
Scan2Cap. As shown in Table 9, we provide a detailed comparison on the validation set of Scan2Cap [6]. During inference, we utilize the object proposals from [18], which include 50 predicted objects extracted with Mask3D [34] for each scan. From the table, we can see our method achieves state-of-the-art results on CIDEr and BLEU-4 at 83.77 and 42.43, respectively.